P($)       Q

70         2800

60         3000


Calculate Price Elasticity of Demand using Mid-Point Method.


1
Expert's answer
2022-01-13T09:01:26-0500

To calculate price elasticity using me point method, you are supposed to use the percentage change in both the quantity and price. The formula is as below;

% change in Quantity"=\frac {Q_2- Q_1}{(Q_2+Q_1)\frac{1}{2}} \times 100"


"=\frac {3000- 2800}{(3000+2800)\frac{1}{2}}\times 100"


"=\frac {200}{2900}\times 100=6.9%"%


% Change in Price"=\frac {P_2- P_1}{(P_2+P_1)\frac{1}{2}}\times 100"

"=\frac {60- 70}{(60+70)\frac{1}{2}}\times 100"


"=\frac {-10}{65}\times 100=-15.4%"%


Price Elasticity "=\frac {6.9}{-15.4}"

= -0.45